OpenStudy (anonymous):

find the sum of the infinite series an= 5-(2/n^2) from n=1 to infinity

OpenStudy (anonymous):

a1 = 3 Sum = n/2(3 + 5 - (2/n²)) sum = n/2(8 - 2/n²) = 4n - 1/n